OpenStudy (anonymous):

a projectile is fired with an initial speed of 500m/s and angle of elevation 30 degrees. find the range, max height and speed at impact

OpenStudy (anonymous):

range = \[12500\sqrt{3}\] max height = 18750 speed at impact= 500m/s

OpenStudy (anonymous):

range = \[2 \times Vox \times t\] with t is the time of projectile on max height position

OpenStudy (anonymous):

the time of projectile on max height position, when Vty (on y axis) is zero So using equation Vt=Voy - at // a = gravity Vt = (Initial speed) x (sin 30) - 10 t 0 = 500 sin30 -10t \[0=250\sqrt{3}-10t\]\[t = 25\sqrt{3}\]

OpenStudy (anonymous):

thanks this helped a lot!